Question One

The manufacturer of a new car claims for it an average mileage of 60 km per litre with a standard deviation of 2 km under city conditions. However, the mileage in 64 trials is found to be 57 km.

Test to determine if the manufacturer’s claim is justified at 5% level of significance.                                                                                                                                                (6 marks)

i)                                  Construct the 99% confidence interval estimate for the mean mileage            (2 marks)



1
Expert's answer
2021-11-29T16:04:29-0500

"H_0:\mu=60" , average mileage is 60 km

"H_a:\mu\neq60" , average mileage is not 60 km


"t=\frac{\overline{x}-\mu}{\sigma/\sqrt n}=\frac{57-60}{2/8}=-12"

"df=n-1=63"

critical value:

"t_{crit}=2"

Since  "|t|>t_{crit}" , we reject the null hypothesis. Average mileage is not 60 km


 for 99% confidence interval:

"t_{crit}=\pm 2.66"


"-2.66<\frac{57-\mu}{2/8}<2.66"


"-0.665<57-\mu<0.665"


"56.335<\mu<57.665"